OpenStudy (anonymous):

The legend on a map shows that 1.5 centimeters equals 60 miles. What distance would be represented by a measurement of 6 centimeters? a. 240 miles c. 200 miles b. 24 miles d. 0.15 mile

OpenStudy (anonymous):

this one is really hard so i need someone to go slow for me lol.

OpenStudy (anonymous):

its another proportion 1.5/60 = 6/x (1.5 cm to 60 miles = 6 cm to x miles) cross multiply (1.5)(x) = (60)(6) 1.5x = 360 can you finish and solve for x (the number of miles) ?

OpenStudy (anonymous):

1.5x/1.5=360/1.5=240

OpenStudy (anonymous):

exactly...you did it...240 miles is correct :)

OpenStudy (anonymous):

thanks(:

OpenStudy (anonymous):

anytime :)